the length {yd.} is 48 the width {yd.} is 1 the area P{yd.2} is 48 what is the perimeter ? A very lost Grandma helping with homework. another is length{yd.} 3 Width{yd.} 16 Area{yd.2} 48 ????
Answer by advanced_Learner(501)

part one
Lenght=48
Width=1
area is =48
perimeter=2lenght+2width=96+2=98
part two
lenght=3
width=16
area is=3*16
area=48

perimeter=2*3+2*16=6+32=38
use same formula.